# The Very Faint X-ray Transient 4XMM J174610.7–290020 at the Galactic center

### v1

@@ Abstract
Context. Very Faint X-ray Transients (VFXTs) are a class of X-ray binary systems that exhibit occasional outbursts with peak X-ray luminosities ($L_{X}\lesssim 10^{36}$ erg s-1) much lower than typical X-ray transients. On 22nd February 2024, during its daily Galactic center monitoring, Swift-XRT detected a VFXT, 7 arcmin from Sgr A, dubbing it Swift J174610–290018.
Aims. We aim to characterize the outburst that occurred in 2024, and a second, distinct outburst in 2025 to understand the nature and accretion flow properties of this new VFXT.
Methods. Swift-XRT light curves are used to constrain the duration of the two events. We carried out X-ray spectral analysis exploiting XMM-Newton and NuSTAR data. We used Chandra and XMM-Newton observations of the last 25 years to constrain the quiescent luminosity of the source and to compare the two most recent outbursts with previous detections of the source.
Results. During the 2024 outburst, which lasted about 50 days, the source reached a luminosity in the 2–10 keV band of $L_{2-10}\simeq 1.2\times 10^{35}$ erg s-1 (assuming it is located at the Galactic center, i.e., at a distance of 8.2 kpc). The 2025 outburst is shorter (about 5 days), and reached $L_{2-10}\simeq 9\times 10^{34}$ erg s-1. The spectral features of the source include an excess at 6.5–7 keV, which can be associated either with a single reflection line or with the ionized Fe XXV and XXVI lines. The same source was identified in both the XMM-Newton and Chandra catalogs of point sources (known as 4XMM J174610.7–290020 and 2CXO J174610.7–290019). During previous detections the source displayed luminosity levels ranging from $L_{2-10}\simeq 2\times 10^{32}$ to $L_{2-10}\simeq 3\times 10^{34}$ erg s-1 between 2000 and 2010. Moreover, it exhibited a potential type I X-ray burst in 2004.
Conclusions. The analysis of the outbursts and the potential type I burst strongly suggests the neutron star low mass X-ray binary (NS-LMXB) nature of the VFXT 4XMM J174610.7–290020. The source can be described by an accretion disk corona (as has been recently proposed by the XRISM/Xtend analysis). This scenario explains the overall low luminosity of this transient and the peculiar iron lines in the spectrum.
@@ 1 Introduction
The Galactic Center (GC), being one of the most crowded and frequently observed regions of the Milky Way, has enabled the discovery of a few tens of X-ray transients (Watson et al., 1981; Pavlinsky et al., 1994; Sidoli et al., 1999, 2001; Sakano et al., 2002; Muno et al., 2003, 2005, 2009; Wijnands et al., 2006; Kuulkers et al., 2007; Ponti et al., 2015; Degenaar et al., 2012, 2015; Mori et al., 2021). Many of those sources are X-ray binaries that spend most of the time in quiescence with X-ray luminosity $L_{X}\lesssim 10^{33}$ erg s-1, but occasionally show outbursts with peaks of $L_{X}\sim 10^{37}-10^{39}$ erg s-1 (Degenaar et al., 2012). A population of hundreds of X-ray binaries is believed to lie in the central parsec of the Milky Way (Muno et al., 2005; Hailey et al., 2018; Mori et al., 2021). Some of the X-ray transients exhibit peak luminosities considerably lower than those of typical X-ray binaries ($L_{X}\lesssim 10^{36}$ erg s-1). This class of transients is usually referred to as Very Faint X-ray Transients (VFXTs, Wijnands et al., 2006, 2015).
@@
About a few tens of VFXTs have been detected in the Galaxy, and about a dozen in the GC (Muno et al., 2005; Degenaar et al., 2012; Bahramian & Degenaar, 2023). The overabundance of these objects at the GC is probably due to an observation bias. Indeed, the GC is the most frequently observed region of the Milky Way in the X-ray band. Almost every year, observations of the GC are performed with Chandra and XMM-Newton (from 2000); NuSTAR (from 2012); and recently with the Einstein Probe (EP). Moreover, since 2006, Swift-XRT performs short ($\sim 1$ ks) observations of the GC almost daily [^1] Due to the Sun position, the GC is not observable between the end of November and the beginning of February., which allows the detection of transients down to $L_{X}\simeq 10^{34}$ erg s-1, in the inner 15 arcmin of the Galaxy (Degenaar et al., 2015).
@@
Given their typical low fluxes, the number of VFXTs that have been studied in detail is still modest; therefore, there is no clear, coherent picture yet of the demographics of this population. While the majority of VFXTs present outbursts that last for a few weeks (Degenaar et al., 2015), a few persistent and quasi-persistent VFXTs have been detected (Bahramian & Degenaar, 2023). The VFXT class is heterogeneous, comprising sources of different natures. While a small fraction of VFXTs can be intrinsically bright, appearing faint due to inclination effects (King & Wijnands, 2006), most VFXTs are thought to be low-mass X-ray binaries (LMXBs) accreting at a low rate, thus explaining their relatively low luminosity (Wijnands et al., 2006; Degenaar et al., 2015). Indeed, almost half of the studied VFXTs have been confirmed as LMXBs in which the compact object is a neutron star (NS-LMXB), via the detection of type I X-ray bursts [^2] Type I bursts are thermonuclear flashes on the surface of an accreting neutron star in a LMXB, caused by unstable burning of accumulated hydrogen and/or helium (Lewin et al., 1993; Strohmayer & Bildsten, 2003). or coherent pulsations at the NS spin period; while in a few cases the binary system involves a black hole (BH-LMXB, Bahramian & Degenaar, 2023).
@@
Various mechanisms have been introduced to explain the relatively low accretion rates of VFXTs. In the NS-LMXB case, a relatively strong enough magnetic field may quench the accretion flow via the magnetic propeller effect (Degenaar et al., 2014, 2017; Heinke et al., 2015; van den Eijnden et al., 2018). Alternatively, the low luminosity may be the result of a small accretion disk. This scenario can be achieved with the compact object accreting from a hydrogen-depleted or planetary companions in very compact orbits (King & Wijnands, 2006; Hameury & Lasota, 2016), such as ultra-compact X-ray binaries (UCXBs). Studying new VFXTs can reveal interesting subtypes of LMXBs, of which only a modest number have been discovered to date – thereby tracing an accretion regime that is still relatively unexplored.
@@
Beyond LMXBs, with an X-ray luminosity $L_{X}\lesssim 10^{36}$ erg s-1, high mass X-ray binaries (HMXBs) can also belong to the VFXT class (Wijnands et al., 2006; Mandel et al., 2025). In this case the compact object is typically accreting from the wind of the companion and an accretion disk might not be present. Moreover, accreting white dwarfs cataclysmic variables (CV) can contribute to the faint end of the VFXT population (Hofmann et al., 2018). Other than binary systems, also magnetars that have typical quiescent luminosity of $L_{X}\sim 10^{33}$ erg s-1 (Mereghetti et al., 2015), can undergo faint outbursts consistent with the VFXTs fluxes (Coti Zelati et al., 2018) and contribute a small fraction of the VFXT population.
@@
In this work, we study the outburst of an X-ray transient that was discovered by the Swift-XRT telescope on 22 February 2024 (Reynolds et al., 2024) during the daily monitoring of the GC (Degenaar et al., 2015). The source, named Swift J174610–290018, was proposed as a new VFXT. The transient is located 6.7 arcmin east of Sgr A (Fig. 1), corresponding to a projected distance of 16 pc at 8.2 kpc (GRAVITY Collaboration et al., 2019), with an initial reported unabsorbed flux of $F_{2-10}=(6.1\pm 2.4)\times 10^{-12}$ erg cm-2 s-1 in the 2–10 keV band, corresponding to a luminosity at the GC distance of $L_{2-10}\simeq 5\times 10^{34}$ erg s-1 (Reynolds et al., 2024). The outburst was also observed with XMM-Newton, NuSTAR, and EP. The Swift-XRT monitoring have shown that the source returned to quiescence at the beginning of April 2024 (Degenaar et al., 2024).
@@
This transient was also observed on 26 February 2024 by XRISM with the XTend instrument (the imaging camera), which shows a peculiar spectrum consistent with a NS-LMXB in addition to ionized iron lines (Fe XXV and Fe XXVI) at $\sim 7$ keV (Yoshimoto et al., 2025). The authors dubbed the source XRISM J174610.8–290021 and proposed it as an accretion disk corona source (White & Holt, 1982; Dove et al., 1997; Iaria et al., 2013), in which the direct emission from the NS-LMXB is blocked by the thick accretion flow, as the system is oriented edge-on. The observed flux is then the radiation scattered by the hot plasma, photoionized by the NS-LMXB, that surrounds the accretion flow (the accretion disk corona).
@@
Grefenstette et al. (2025) reported a rebrightening of the same source, serendipitously observed with NuSTAR, on 4 April 2025, with a spectrum and flux similar to those of the 2024 outburst. The source remained detectable in subsequent NuSTAR observations over the few following days, before returning to quiescence (Degenaar et al., 2025).
@@
The article is organized as follows: in Sect. 2, we illustrate the dataset used in this study. In Sect. 3, we characterize the outburst that occurred in 2024. Section 4 describes the outburst observed in 2025, while an archival search revealing previous detections of the source is addressed in Sect. 5. In Sect. 6 we discuss the observational results.
@@
Figure 1: EPIC-pn image in the 2–10 keV band of the GC, obtained from the 5 XMM-Newton observations taken in the Spring of 2024. Swift J174610–290018 is located 6.7 arcmin east of Sgr A. The eclipsing NS-LMXB AX J1745–2901 (Maeda et al., 1996) was active during the XMM-Newton observations and it is the brightest source in the field. The inset displays the position uncertainties ($1\sigma$) of the 2024 transient (green circle), and the known X-ray source reported in the 4XMM-DR14 and CSC 2.1 catalogs.
@@ 2 X-ray observations
Table 1: Log of the observations used for spectral analysis in this article. For XMM-Newton, the exposure reported in parentheses is the one after filtering for flaring particle background.
@@ 2.1 Swift
We use Swift-XRT data of the GC monitoring campaign (Degenaar et al., 2015) for creating the light curves of the 2024 and 2025 outbursts. To this end, we used obsIDs 00096991167–00097302020 performed between 2024-02-09 and 2024-04-21, and obsIDs 00097302236–00097302245 performed between 2025-04-02 and 2025-04-11. All observations were taken in Photon Counting (PC) mode and exposure times ranged from 0.70 to 1.21 ks. The raw Swift-XRT data were processed with the xrtpipeline using standard quality cuts and event grades 0–12. Light curves were created by extracting source events with XSelect using a 18-arcsecond circular region centered on the source position.
@@ 2.2 XMM-Newton
The outburst of the source is detected in five observations of the GC taken in the Spring of 2024 by XMM-Newton (Jansen et al., 2001). We reprocessed the raw data using the standard pipelines in the Scientific Analysis System of XMM-Newton (SAS 21.0), using the last release of the current calibration files (CCF) as of April 2024. Given the higher effective area of EPIC-pn (Strüder et al., 2001) compared to the MOS (Turner et al., 2001), we used only the first camera in our analysis. We filtered the event list for high particle background by creating light curves of the entire CCD, in the 10–12 keV band, with a 100 s time bin. We then defined Good-Time-Intervals (GTI) as the periods in which the count rate is below 0.4 cts s-1.
@@ 2.2.1 Source position
Given the source’s count rate, its relative position on the CCD can be determined with a statistical uncertainty of approximately 0.1–0.2 arcsec. However, the overall positional uncertainty is dominated by the systematic error in XMM-Newton absolute astrometry. To account for this, we retrieved the astrometrically corrected source coordinates from the Pipeline Processing System (PPS) products for each observation, and performed a fit to determine the RA, Dec, and associated positional uncertainty, assuming Gaussian errors. The best position of the source is RA, Dec = (17h46m10.68s, -2900’20.1”) with a 0.7 arcsec uncertainty ($1\sigma$). The position is fully consistent with the source 4XMM J174610.7–290020, as reported in the XMM-Newton Serendipitous Source Catalogue (4XMM-DR14, Webb et al., 2020). In Fig. 1, we show the $1\sigma$ uncertainty of the source analyzed in this work (green circle) and the XMM-Newton source reported in the catalog. 4XMM J174610.7–290020 was previously detected in four XMM-Newton observations taken in 2000 and 2004. Given the consistent position, the fact that both sources are heavily absorbed and show similar spectral lines (Pastor-Marazuela et al., 2020), it is likely that they are the same source, observed in two different intensity states. For the remainder of this work, we consider them to be the same source.
@@ 2.2.2 NIR counterpart
We searched for a Near-Infrared (NIR) counterpart in the GALACTICNUCLEUS catalog (Nogueras-Lara et al., 2021). Within 1.5 arcsec of the best source position reported before, there are 4 NIR sources (we excluded one foreground) from the GALACTICNUCLEUS catalog. Table 2 reports the separation of the sources, their $K_{s}$ band magnitude, and, when available, the $H$ band magnitude.
@@
Table 2: Potential NIR counterparts from the GALACTICNUCLEUS catalog (Nogueras-Lara et al., 2021). $\delta$ is the angular separation between the NIR source and 4XMM J174610.7–290020. For all the NIR sources the uncertainty on RA and Dec is of the mas order, while the error on the magnitudes is typically 0.01 mag.
Given the limitations of current NIR surveys, which are largely incomplete at $K_{s}\sim 16$th mag due to crowding in the dense GC region, detecting a non-evolved low-mass donor star is highly improbable. We therefore consider it likely that the neither of the four NIR sources listed above are the true NIR counterpart to 4XMM J174610.7–290020.

### v2

# The very faint X-ray transient Swift J174610-290018 at the Galactic center

@@ Abstract
Context. Very faint X-ray transients (VFXTs) are a class of X-ray binary systems that exhibit occasional outbursts with peak X-ray luminosities ($L_{X}\lesssim 10^{36}$ erg s-1) much lower than those of typical X-ray transients. On 22 February 2024, during its daily Galactic center monitoring, Swift-XRT detected a VFXT 7 arcmin from Sgr A, dubbing it Swift J174610-290018.
Aims. We aim to characterize the outburst that occurred in 2024 and a second, distinct outburst in 2025 to understand the nature and accretion flow properties of this new VFXT.
Methods. Swift-XRT light curves were used to constrain the duration of the two events. We carried out X-ray spectral analysis, exploiting XMM-Newton and NuSTAR data. We used Chandra and XMM-Newton observations from the last 25 years to constrain the quiescent luminosity of the source and to compare the two most recent outbursts with previous detections of the source.
Results. During the 2024 outburst, which lasted about 50 days, the source reached a luminosity in the 2–10 keV band ($L_{2-10}$) of $\simeq 1.2\times 10^{35}$ erg s-1 (assuming it is located at the Galactic center, i.e., at a distance of 8.2 kpc). The 2025 outburst is shorter (about 5 days) and reached $L_{2-10}\simeq 9\times 10^{34}$ erg s-1. The spectral features of the source include an excess at 6.5–7 keV, which can be associated either with a single reflection line or with the ionized Fe XXV and XXVI lines. The same source was identified in both the XMM-Newton and Chandra catalogs of point sources (known as 4XMM J174610.7-290020 and 2CXO J174610.7-290019). During previous detections the source displayed luminosity levels ranging from $L_{2-10}\simeq 2\times 10^{32}$ to $L_{2-10}\simeq 3\times 10^{33}$ erg s-1 between 2000 and 2010. Moreover, it exhibited a potential type I X-ray burst in 2004.
Conclusions. The analysis of the outbursts and the potential type I burst strongly suggests that the VFXT Swift J174610-290018 is a neutron star low-mass X-ray binary. The source can be described as being an accretion disk corona source (as has been recently proposed by an XRISM/Xtend analysis). This scenario explains the overall low luminosity of this transient and the peculiar iron lines in the spectrum.
@@ 1 Introduction
The Galactic center (GC), being one of the most crowded and frequently observed regions of the Milky Way, has enabled the discovery of a few dozen X-ray transients (Watson1981; Pavlinsky1994; Sidoli1999; Sidoli2001; Sakano2002; Muno2003; Muno2005; Muno2009; Wijnands2006; Kuulkers2007; Ponti2015; Degenaar2012; Degenaar2015; Mori2021). Many of those sources are X-ray binaries that spend most of their time in quiescence with X-ray luminosities ($L_{X}$) of $\lesssim 10^{33}$ erg s-1 but occasionally show outbursts with peaks of $\sim 10^{37}-10^{39}$ erg s-1 (Degenaar2012). A population of hundreds of X-ray binaries is believed to lie in the central parsec of the Milky Way (Muno2005; Hailey2018; Mori2021). Some of the X-ray transients exhibit peak luminosities considerably lower than those of typical X-ray binaries ($L_{X}\lesssim 10^{36}$ erg s-1). These are usually referred to as very faint X-ray transients (VFXTs; Wijnands2006; Wijnands2015).
@@
A few dozen VFXTs have been detected in the Galaxy, and about a dozen in the GC (Muno2005; Degenaar2012; Bahramian2023). The overabundance of these objects at the GC is probably due to an observation bias. Indeed, the GC is the most frequently observed region of the Milky Way in the X-ray band. Almost every year, observations of the GC are performed with Chandra and XMM-Newton (from 2000); NuSTAR (from 2012); and recently with the Einstein Probe (EP). Moreover, since 2006, the Swift X-Ray Telescope (Swift-XRT) has performed short ($\sim 1$ ks) observations of the GC almost daily [^1] Due to the Sun position, the GC is not observable between the end of November and the beginning of February., which allows the detection of transients down to $L_{X}\simeq 10^{34}$ erg s-1 in the inner 15 arcmin of the Galaxy (Degenaar2015).
@@
Given their typical low fluxes, the number of VFXTs that have been studied in detail is still modest; therefore, there is no clear, coherent picture yet of the demographics of this population. While the majority of VFXTs present outbursts that last for a few weeks (Degenaar2015), a few persistent and quasi-persistent VFXTs have been detected (Bahramian2023). The VFXT class is heterogeneous, comprising sources of different natures. While a small fraction of VFXTs can be intrinsically bright, appearing faint due to inclination effects (King2006), most VFXTs are thought to be low-mass X-ray binaries (LMXBs) accreting at a low rate, thus explaining their relatively low luminosity (Wijnands2006; Degenaar2015). Almost half of the studied VFXTs have been confirmed as LMXBs in which the compact object is a neutron star (NS) via the detection of type I X-ray bursts [^2] Type I bursts are thermonuclear flashes on the surface of an accreting NS in a LMXB, caused by unstable burning of accumulated hydrogen and/or helium (Lewin1993; Strohmayer2003). or coherent pulsations at the NS spin period; in a few cases, the binary system involves a black hole (Bahramian2023).
@@
Various mechanisms have been introduced to explain the relatively low accretion rates of VFXTs. In the NS-LMXB case, a sufficiently strong magnetic field may quench the accretion flow via the magnetic propeller effect (Degenaar2014; Degenaar2017; Heinke2015; Eijnden2018). Alternatively, the low luminosity may be the result of a small accretion disk. This scenario would occur if the compact object accretes from a hydrogen-depleted or planetary companion in very compact orbits (King2006; Hameury2016), such as ultra-compact X-ray binaries. Studying new VFXTs can reveal interesting subtypes of LMXBs, of which only a modest number have been discovered to date – thereby tracing an accretion regime that is still relatively unexplored.
@@
High-mass X-ray binaries (HMXBs) with X-ray luminosities ($L_{X}$) of $\lesssim 10^{36}$ erg s-1, can also belong to the VFXT class (Wijnands2006; Mandel2025). In this case, the compact object is typically accreting from the wind of the companion and an accretion disk might not be present. Moreover, accreting white dwarf cataclysmic variables (CVs) can contribute to the faint end of the VFXT population (Hofmann2018). In addition to binary systems, magnetars that have typical quiescent luminosities ($L_{X}$) of $\sim 10^{33}$ erg s-1 (Mereghetti2015) can also undergo faint outbursts consistent with VFXTs fluxes (Coti_Zelati2018); they make up a small fraction of the VFXT population.
@@
In this work we studied the outburst of an X-ray transient that was discovered by the Swift-XRT telescope on 22 February 2024 (Reynolds2024) during the daily monitoring of the GC (Degenaar2015). The source, named Swift J174610-290018, was proposed as a new VFXT. The transient is located 6.7 arcmin east of Sgr A (Fig. 1), corresponding to a projected distance of 16 pc at 8.2 kpc (Gravity2019), with an initial reported unabsorbed flux in the 2–10 keV band ($F_{2-10}$) of $(6.1\pm 2.4)\times 10^{-12}$ erg cm-2 s-1, corresponding to a luminosity ($L_{2-10}$) of $\simeq 5\times 10^{34}$ erg s-1 at the GC distance (Reynolds2024). The outburst was also observed with XMM-Newton, NuSTAR, and EP. Swift-XRT monitoring has shown that the source returned to quiescence at the beginning of April 2024 (Degenaar2024).
@@
This transient was also observed on 26 February 2024 by the X-Ray Imaging and Spectroscopy Mission (XRISM) with the XTend instrument (the imaging camera); it showed a peculiar spectrum consistent with a NS-LMXB in addition to ionized iron lines (Fe XXV and Fe XXVI) at $\sim 7$ keV (Yoshimoto2025). The authors dubbed the source XRISM J174610.8-290021 and proposed it to be an accretion disk corona source (White1982; Dove1997; Iaria2013) in which the direct emission from the NS-LMXB is blocked by the thick accretion flow, as the system is oriented edge-on. The observed flux is then the radiation-scattered by the ionized plasma that surrounds the accretion flow (the accretion disk corona).
@@
Grefenstette2025 report a re-brightening of the same source, serendipitously observed with NuSTAR on 4 April 2025, with a spectrum and flux similar to those of the 2024 outburst. The source remained detectable in subsequent NuSTAR observations over the following few days, before returning to quiescence (Degenaar2025).
@@
The article is organized as follows: In Sect. 2 we illustrate the dataset used in this study. In Sect. 3 we characterize the outburst that occurred in 2024. Section 4 describes the outburst observed in 2025, and an archival search revealing previous detections of the source is addressed in Sect. 5. In Sect. 6 we discuss the observational results.
@@
Figure 1: EPIC-pn image in the 2–10 keV band of the GC, obtained from the five XMM-Newton observations taken in the spring of 2024. Swift J174610-290018 is located 6.7 arcmin east of Sgr A. The eclipsing NS-LMXB AX J1745–2901 (Maeda1996) was active during the XMM-Newton observations and is the brightest source in the field. The inset displays the position uncertainties ($1\sigma$) of the 2024 transient (green circle) and the known X-ray source reported in the 4XMM-DR14 and CSC 2.1 catalogs.
@@ 2 X-ray observations
Table 1: Log of the observations used for spectral analysis in this article. [^3] For XMM-Newton, the exposure reported in parentheses is the one after filtering for flaring particle background.
@@ 2.1 Swift
We used Swift-XRT data of the GC monitoring campaign (Degenaar2015) to create the light curves of the 2024 and 2025 outbursts. To this end, we used obsIDs 00096991167–00097302020 performed between 9 February and 21 April 2024, and obsIDs 00097302236–00097302245 performed between 2 and 11 April 2025. All observations were taken in photon counting mode and exposure times ranged from 0.70 to 1.21 ks. The raw Swift-XRT data were processed with the xrtpipeline using standard quality cuts and event grades 0–12. Light curves were created by extracting source events with XSelect using a 18-arcsecond circular region centered on the source position.
@@ 2.2 XMM-Newton
The outburst of the source is detected in five observations of the GC taken in the Spring of 2024 by XMM-Newton (Jansen2001). We reprocessed the raw data using the standard pipelines in the Scientific Analysis System of XMM-Newton (SAS 21.0), using the last release of the current calibration files as of April 2024. Given the higher effective area of EPIC-pn (Struder2001) compared to the MOS (Turner2001), we used only the first camera in our analysis. We filtered the event list for high particle background by creating light curves of the entire CCD, in the 10–12 keV band, with a 100 s time bin. We then defined good time intervals as the periods in which the count rate is below 0.4 cts s-1.
@@ 2.2.1 Source position
Given the source’s count rate, its relative position on the CCD can be determined with a statistical uncertainty of approximately 0.1–0.2 arcsec. However, the overall positional uncertainty is dominated by the systematic error in XMM-Newton absolute astrometry. To account for this, we retrieved the astrometrically corrected source coordinates from the Pipeline Processing System products for each observation, and performed a fit to determine the RA, Dec, and associated positional uncertainty, assuming Gaussian errors. The best position of the source is RA, Dec = (17h46m10.68s, -2900’20.1”) with a 0.7 arcsec uncertainty ($1\sigma$). The position is fully consistent with the source 4XMM J174610.7-290020, as reported in the XMM-Newton Serendipitous Source Catalogue (4XMM-DR14; Webb2020). In Fig. 1 we show the $1\sigma$ uncertainty of the source analyzed in this work (green circle) and the XMM-Newton source reported in the catalog. 4XMM J174610.7-290020 was previously detected in four XMM-Newton observations taken in 2000 and 2004. Given the consistent position, the fact that both sources are heavily absorbed and show similar spectral lines (Pastor-Marazuela2020), it is likely that they are the same source, observed in two different intensity states. For the remainder of this work, we consider them to be the same source.
@@ 2.2.2 NIR counterpart
We searched for a near-infrared (NIR) counterpart in the GALACTICNUCLEUS catalog (Nogueras2021). Within 1.5 arcsec of the best source position reported before, there are 4 NIR sources (we excluded one foreground) from the GALACTICNUCLEUS catalog. Table 5 reports the separation of the sources, their $K_{s}$ band magnitude, and, when available, the $H$ band magnitude.
[…]
@@
Given the limitations of current NIR surveys, which are largely incomplete at $K_{s}\sim 16$th mag due to crowding in the dense GC region, detecting an unevolved low-mass donor star is highly improbable. We therefore consider it likely that none of the four NIR sources listed above is the true NIR counterpart to Swift J174610-290018.